OpenStudy (anonymous):

The East-Side soccer team has a cylindrical water cooler that is half full. The diameter of the cooler is 12 in. and the height is 15 in. The team’s drinking cups hold 20.14 in3. How many drinking cups can be filled from the cooler? Use 3.14 to approximate pi, and round your answer to the nearest whole number. A. 14 B. 42 C. 84 D. 169

OpenStudy (anonymous):

i think its b?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Nnesha do u no it?

OpenStudy (anonymous):

The equation for the volume (which we need to find, since we're told the cups are filled in inch^3) of a cylinder is: \[\pi*r ^{2}*h\]We know this because a cyclinder is just a circle that is extended by h height, so think of it as the area of a circel that is multiplied (or extended) by h. We're also told that the cooler is half full, so we divide that equation by 2. So now we need to find the total volume of liquid in the cooler. Just sub in the height and half the diameter (radius = 6 inch). I got 848 inch^3 for the juice volume. Now divide that by the volume in a cup (20.42in3) You should get 42, which is B.

OpenStudy (anonymous):

A jar of mayonnaise is shaped like a cylinder. The radius of the jar is 1.5 in. and its height is 7 in. The mayonnaise is on sale for $0.08 per in3. What is the cost of the jar of mayonnaise, rounded to the nearest cent? Use 3.14 to approximate pi. $_____

OpenStudy (anonymous):

do you no this? @iGreen

OpenStudy (anonymous):

This is the same application of the equation of a volume of a cylinder that was mentioned. Just sub in the radius and the height, then multiply the entire equation by 0.08.

OpenStudy (anonymous):

so is it 3.26? @Hari_Seldon

OpenStudy (anonymous):

No, it's $3.96. pi*(1.5)^2 * 7 = 49.48 in3 49.48 * 0.08 = 3.958 = $3.96

OpenStudy (anonymous):

oooh ok can u help me with 2 more? @Hari_Seldon